What is the RPM of a Verticle [sic] Wheel if we have following things:
1) Diameter : 6 mtrs
2) Velocity : 130 m/s
3) Volume : 20.4 ltrs/s
4) Flow : 10 Deg to TOP
5) Flow Dis.: 1 mtr above wheel
5) RPM : ?

Answer by Alan3354(69443)

What is the RPM of a Verticle Wheel if we have following things:
1) Diameter : 6 mtrs
2) Velocity : 130 m/s
-----------------------
130 m/sec = 7800 m/minute
---
Rev/min = 7800/(6*pi) = 1300/pi
RPM =~ 413.8